OpenStudy (anonymous):

in a right triangle the length of a hypotenuse is c and the length of one leg is a and and the other leg is b, what is the value of b if: c=2b, a=12

OpenStudy (debbieg):

Use the Pythagorean Theorem: \(\large a^2+b^2=c^2\) You have: a=12, and c=2b So: \(\large 12^2+b^2=(2b)^2\) Solve that for b. The answer is NOT b=12.

OpenStudy (debbieg):

OpenStudy values the Learning process - not the ‘Give you an answer’ process. Don’t post only answers - guide the asker to a solution. http://openstudy.com/code-of-conduct Please dont post answers @scofflawxx. Especially wrong ones (but thank you for deleting it).